An electric ceiling fan is rotating about a fixed axis with an initial angular velocity magnitude of 0.200 rev/s . The magnitude of the angular acceleration is 0.886 rev . Both the angular velocity and angular acceleration are directed clockwise. The electric ceiling fan blades form a circle of diameter 0.770m. A). Compute the fans angular velocity magnitude after time 0.198 s has passed. B). Through how many revolutions has the blade turned in the time interval 0.198 s from Part A? C). What is the tangential speed Vianlt) of a point on the tip of the blade at time t 0.198 s? 2. Yo-Yo man releases a yo-yo from rest and allows it to drop, as he keeps the top end of the string stationary. The mass of the yo-yo is 0.056 kg, its moment of inertia is 2.9x10-Skg m2 and the radius, r, of the axle the string wraps around is 0.0064 m Through what height must the yo-yo fall for its linear speed to be 0.75m/s?
Answer #1

1)

a)

W = R1 + α*t

W = 0.200 rev/s + (0.886 rev/s^2 x 0.198 s)

W = 0.375 rev/s

b)

Θ = R1*t + 1/2α*t^2

Θ = 0.200*0.198 + 1/2*0.886*0.198^2

Θ = 0.0570 rev

C)

Vθ = rω

Vθ(0.198) = (0.385 m)(0.375 rev/s)(2π rad/rev)

Vθ(0.198) = 0.91 m/s